# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ D.4. Point Cloud Corruption due to Occlusion
    Apply Algorithm 2 (or Algorithm 4) with inputs $\{(\widetilde{\mathbf{X}^{s}},\widetilde{\mathrm{p}^{s}})\}_{s\in[S]}$, $(\widetilde{\mathbf{Y}},\widetilde{\mathrm{q}})$ to compute estimated coordinates $\widetilde{\lambda}$. After that, apply Algorithm 1 (or Algorithm 4) with inputs $\{({\mathbf{X}^{s}},{\mathrm{p}^{s}})\}_{s\in[S]}$, $\widetilde{\lambda}$, ${\mathrm{q}}$, $M$ in order to synthesize a barycenter $\mathbf{Y}_{\widetilde{\lambda}}$ from the original (unperturbed) templates. We refer to $\mathbf{Y}_{\widetilde{\lambda}}$ as a reconstruction of $\mathbf{Y}$. Apply MDS embedding (with 2 or 3 components) to $\mathbf{Y}_{\widetilde{\lambda}}$ in order to visualize the recovered point cloud.
@@
Figure 22. Occlusion with a circular mask (rotation-invariant) and reconstruction (two examples). Top row: Templates. Bottom row from left to right: Synthetic GW barycenter; its perturbation (via occlusion of the central region using a circular mask); the reconstructed point cloud using our proposed method (i.e., GW barycenter coordinates to the corrupted templates are computed using Algorithm 2 and then used to synthesize a GW barycenter with the unperturbed templates); and a random reconstruction (i.e., a synthesized barycenter using a random weight vector $\lambda\in\Delta_{2}$ and the unperturbed templates).
@@
In the examples shown in Figure 22, the BCM paradigm proves to be an effective approach to image reconstruction, as evidenced by its ability to recover the missing regions of a perturbed Gromov–Wasserstein (GW) barycenter. A synthetic barycenter is partially occluded using a circular mask, simulating structured data loss. Leveraging geometric information from the original templates (top panel), the BCM framework successfully reconstructs the barycenter (bottom right).
Appendix E Wasserstein Barycenters and Gromov-Wasserstein Barycenters
@@
$\mathcal{P}_{2}(\mathbb{R}^{d})\ni\nu\longmapsto\frac{1}{2}\sum_{s=1}^{S}\lambda_{s}\,W^{2}_{2}(\nu,\mu^{s}).$ (63)
@@
$\sum_{s=1}^{S}\lambda_{s}(Id-T^{\nu_{\lambda}\to\mu^{s}})=0\quad\text{almost everywhere w.r.t. }\nu_{\lambda}.$ (65)
Moreover, it can be proven that under certain assumptions critical points (or Karcher means), that is, measures $\nu_{\lambda}$ satisfying (65) and Wasserstein barycenters (i.e., minimizer of the functional described by (63)) coincide (see [68] and [43, Thm. 3.1.15]). As a consequence, to find such a Wasserstein barycenter $\nu_{\lambda}$, one can adopt the following fixed-point scheme based on (Wasserstein) gradient descent on the variance functional (63):
@@
In the context of this paper, that is, under the Gromov–Wasserstein (GW) structure, the objective functional $G_{\lambda}$ (see (31)) plays the role analogous to that of (63) in the Wasserstein setting. Its gradient can be computed using the blow-up technique (see Proposition 4.3), and it serves as the counterpart to the Fréchet derivative in (64). In [14], the authors propose an iterative Fréchet mean algorithm based on the work of Pennec [45], interpreting it as a gradient descent method for minimizing the Fréchet functional $G_{\lambda}$ over the space of finite networks. This approach parallels the procedure described in the Wasserstein setting but is adapted to the GW framework, following Pennec’s ideas on averaging in finite-dimensional Riemannian manifolds. However, to the best of our knowledge, there are no theoretical guarantees in the literature that critical points of $G_{\lambda}$ are minimizers of (4), that is, true GW barycenters. Our Lemma 4.5 represents a step in this direction.

### v2

@@ D.2. Point Cloud Corruption due to Occlusion
[…]
@@
Figure 14. Occlusion with a circular mask (rotation-invariant) and reconstruction (two examples). Top row: Templates. Bottom row from left to right: Synthetic GW barycenter; its perturbation (via occlusion of the central region using a circular mask); the reconstructed point cloud using our proposed method (i.e., GW barycenter coordinates to the corrupted templates are computed using Algorithm 2 and then used to synthesize a GW barycenter with the unperturbed templates); and a random reconstruction (i.e., a synthesized barycenter using a random weight vector $\lambda\in\Delta_{1}$ and the unperturbed templates).
@@
In the examples shown in Figure 14, the BCM paradigm proves to be an effective approach to image reconstruction, as evidenced by its ability to recover the missing regions of a perturbed Gromov-Wasserstein (GW) barycenter. A synthetic barycenter is partially occluded using a circular mask, simulating structured data loss. Leveraging geometric information from the original templates (top panel), the BCM framework successfully reconstructs the barycenter (bottom right).
D.3. Exploring Alternative Node Dissimilarities: Euclidean and Non-Euclidean
From the experimental perspective, further explorations using different node dissimilarities would be interesting to analyze. In this section we show some initial empirical results.
a Templates.
b Target and its reconstruction.
c Noisy target and its reconstruction.
Figure 15. Ambient distance is considered to compute distance matrices between the points of the above shapes, i.e., the Euclidean distance in $\mathbb{R}^{2}$. In (b) and (c), reconstruction is performed by first solving the analysis problem to estimate GW $\lambda$-coordinates with respect to the templates in (a); then synthesizing the corresponding GW barycenter with those weights and templates from (a); and finally using MDS on the recovered matrix for visualization. In (b) the recovered weight via the analysis algorithm 2 is (0.6,0.4).
We consider two fixed templates: a circle and a square shape as illustrated in Fig. 15a. We want to explore how the analysis-synthesis pipeline behaves with different node dissimilarities. Using the natural ambient-space node dissimilarity (Euclidean distance in $\mathbb{R}^{2}$), we observe in Fig. 15b “artifacts” or “corners” that prevent perfect recovery of a circular input (elliptic contour), likely due to the interaction between the square template and the use of the Euclidean metric in the plane. Here, recovery is done by first estimating weights via GW analysis and then synthesizing a GW barycenter with such weights. Nevertheless, when the input ellipse is perturbed by noise, the reconstruction remains reasonably stable (Fig. 15c) in comparison with the original reconstruction with no noise (Fig. 15b).
It is natural to ask ourselves how this would be influenced if intrinsic distances are considered, such as geodesic distances on manifolds. As a preliminary exploration, in Table 16, we consider $p$-Fermat distances for different input shapes (also called “power-weighted shortest-path distances”), see [35, Definition 1.1]. The inputs (or targets) are defined as the set of points $(x,y)\in\mathbb{R}^{2}$ satisfying $|x|^{q}+|y|^{q}=1$ (i.e., boundaries of a $q$-ball in the plane), for different values of $q$. Thus, when $q=2$ the input shape is a circle and when $q=1$, a square. We sample them uniformly.
For the geodesic approach on the templates, each template is represented by an intrinsic distance matrix: for the circle we use the exact arc-length (geodesic) distance along the curve, and for the square we use shortest-path distances along the ordered boundary cycle (i.e., sums of edge lengths along the contour). Moreover, to make the square case more informative (and avoid it effectively reducing to the circle case), we sample the square more densely near its corners while keeping a uniform measure on the sampled points; this concentrates more mass at the corners and accentuates their geometric influence. These choices replace ambient Euclidean distances between node points with distances measured along the shape itself, i.e., within the template’s one-dimensional geometry. In this preliminary exposition, we consider uniform mass across sample nodes in the target and template 1D-shapes as well as equal number of nodes.
Table 16 reports the estimated GW coordinates $\lambda=(\lambda_{1},\lambda_{2})$ for different values of $p$ and $q$ along with the GW distance between the input shape ($q$-boundary ball with its corresponding $p$-Fermat distance matrix) and the recovered one. Reconstruction is done in the same fashion as in Fig. 15: after weight ($\lambda$) estimation though GW-analysis, GW-synthesis is performed. The $\lambda_{1}$ coordinate corresponds to the circle shape and $\lambda_{2}$ to the square template.
The only case in Table 16 where we obtain perfect reconstruction is for $p=1$ and $q=2$. This corresponds to the circle as the target, with its distance matrix computed via a path (shortest-path) metric, since $p=1$ yields the standard graph/path distance. As expected, the reconstruction is essentially exact in this setting $GW\approx 0$ and $\lambda$ concentrated in the first component (circle template). Another notable trend is that for $q=1$ (the square) and for large $q$ (i.e., shapes increasingly close to a square), the barycentric coordinates concentrate on the square template, with $\lambda\approx(0,1)$.
Figure 16. Report of $\lambda$-estimation for target shapes $|x|^{q}+|y|^{q}=1$ with distance matrices calculated via $p$-Fermat distances [35, Definition 1.1], for different values of $q$ (columns) and $p$ (rows). The templates in Fig. 15a are considered with internal geodesic distances. Shape reconstruction is performed by solving first the GW-analysis problem (via Algorithm 2), and then synthesizing with the estimated weights and the fixed templates for all the experiments. GW distances between the initial target and its reconstruction are presented for each $(p,q)$-cell.
Appendix E The Analysis Problem Outside the Barycenter Span: A Preliminary and Empirical Exploration
Assume that $\{\mathbb{X}^{s}\}_{s\in[S]}$ are fixed templates, and recall that we denote the elements of their GW barycenter span as $\mathbb{Y}_{\lambda}$, for $\lambda\in\Delta_{S-1}$. The objective of this section is to visualize, for an arbitrary target $\mathbb{Y}$, the heat map of the distance functional
(63) $\mathbf{J}(\lambda):=GW^{2}(\mathbb{Y}_{\lambda},\mathbb{Y}),$
on the simplex $\Delta_{S-1}$, and the relation between its minima and the output of the analysis Algorithms 2 and 4 when they are applied to the target $\mathbb{Y}$.
By showing how such functional ‘looks like’ over the simplex of weights, we aim to provide preliminary experimental results to the following questions: How the proposed GW analysis methods perform when the target shape $\mathbb{Y}$ does not belong to the GW barycenter space spanned by predetermined templates? What happens when the minimum of the above functional (63) is non-zero or cannot be attained exactly? Do the weights $\lambda$ produced by the proposed GW analysis algorithms coincide with, or at least lie near, a minimizer of $\mathbf{J}$?
Heatmap Generation
To visualize the heatmaps of $\mathbf{J}$ we will have to:
  1. (1)
    Generate templates and a target.
  2. (2)
    Create a grid to discretize the simplex $\Delta_{S-1}$ and synthesize one GW barycenter per point $\lambda$; and
  3. (3)
    Compute the GW distance between the target and each synthesized barycenter of the grid.
Assume that we are given three fixed templates ($S=3$) and a target. If we naively follow the steps above and synthesize GW barycenters in step 2 using already available libraries like ot.gromov [22], when running the procedure again we encounter different landscapes as seen in Fig. 17. This behavior is mostly due to the issues inherent to instability of the GW barycenter synthesis problem. Although improving the synthesis algorithms is not essentially in the scope of this work, we added some tests to mitigate this variability.
Figure 17. Heatmaps illustrating the landscape of $\mathbf{J}$ (63) where no safeguards for GW barycenter synthesis process are applied for obtaining $\{\mathbb{Y}_{\lambda}\}_{\lambda\in\Delta_{2}}$. For the three plots, the same three fixed templates, and the same target $\mathbb{Y}$. Variability across the three landscapes is evident, even though the selected target lies within the GW barycenter space, as indicated by the red cross (weight).
Safeguards for the GW Barycenter Synthesis.
The synthesis procedure requires to solve the non-convex minimization problem
(64) $\mathbb{Y}_{\lambda}\in\operatorname*{arg\,min}_{\mathbb{Z}\in\mathrm{Bary}(\{[\mathbb{X}^{s}]\}_{s\in[S]})}\sum_{s=1}^{S}\lambda_{s}GW^{2}(\mathbb{Z},\mathbb{X}^{s}).$
Non-convex problems often have no guarantees of uniqueness for global minimizers (for the same weight $\lambda$), and moreover, numerical solvers might obtain only local minima or may suffer from convergence issues. In particular, the function ot.gromov.gromov_barycenters from the POT Library [22], used to generate GW barycenters turns to be numerically sensitive under different initializations: if we fix a size $M$, a vector $\mathrm{q}$ of weights, a set of templates $\{\mathbb{X}^{s}\}_{s\in[S]}$, and $\lambda\in\Delta_{S-1}$ and then use such function twice to synthesize $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$ (with the same inputs), they do not necessarily have GW distance $\approx 0$.
This already indicates that the landscapes for $\mathbf{J}$ (with the same target $\mathbb{Y}$) are going to change. Even worse, it also indicates that the synthesized object might not even be a GW barycenter. We added two safeguards to make sure that that is not the case:
  1. (a)
    After synthesizing with the POT Library function, we test if the result is a weak GW barycenter by following Remark 4.9.
  2. (b)
    For a fixed $\lambda$ we generate several candidate barycenters, and choose the one that has minimum synthesis cost (i.e., minimum for the functional in (64)).
With this in mind, we created templates and targets inside and outside of the GW barycenter space, and reproduced the landscapes in Fig. 19 and Fig. 20, respectively.
Template Generation.
We considered $S=3$ and generated three random point clouds of $N$ points each. For each, we assign the Euclidean distance between points as distance matrix, and assumed a uniform measure. In practice, using uniform measures helps make the POT Library’s barycenter synthesis solver output compatible with the blow-up procedure (as empirically noted in Figures 6 and 7), which in turn helps satisfy safeguard (a). We additionally tested the pairwise GW distance between templates to be bigger than a minimum threshold to ensure that the distance between grid points would end being (heuristically) bigger than the distance between different synthesized barycenter with the same weights. See Fig. 18 for a visualization of a choice of such type of templates.
Figure 18. Orange, blue and green point clouds illustrating a choice of 3 templates.
Target Generation.
(1) To obtain a target outside the barycenter space, we generated one extra random point cloud of size $N$ as when creating the templates. (2) To obtain a target in the GW barycenter space of the templates, we artificially synthesized a new GW barycenter with a chosen weight $\lambda$ using the safeguards mentioned above. In our visualizations we considered $\lambda=(0.3,0.4,0.3)$. Since the weight $\lambda$ was taken inside the grid, we made sure that the target was created before (and independently) of the grid barycenters to avoid forcing a zero of $\mathbf{J}$ at that point.
Conclusions and Observations.
Through this brute-force method for visualizing the $\mathbf{J}$ landscapes (Figures 19 and 20), one can informally assess the relation between the proposed analysis Algorithms 2 and 4 and the minima of the functional $\mathbf{J}$ for targets inside and outside the GW barycenter space. When the target lies in the barycenter space, both algorithms recover the expected weights. When outside the GW barycenter space, the recovered weights still coincide for both algorithms. In both cases, the recovered weights seem to lie in regions where $\mathbf{J}$ is small (deep blue regions in the heatmaps) showing qualitative agreement with the landscape structure. Consequently, the out-of-model results are satisfactory.
It is worth noticing that the lack of uniqueness for the synthesis problem (from both, theoretical and computational, perspectives) makes it very difficult to properly visualize the global minima of $\mathbf{J}$ for targets outside the barycenter space. Even though we included all the technicalities mentioned above, trying to guarantee less variability of the (brute-force) heatmaps, still the three plots within Figures 19 and 20, respectively, do not coincide exactly.
Figure 19. Heatmaps illustrating the landscape of $\mathbf{J}$ over the first two components of the simplex $\Delta_{2}$. For the same three fixed templates and the same target $\mathbb{Y}$ lying in their GW barycenter space, we evaluate $\mathbf{J}(\lambda^{(i)})$ on a grid $\{\lambda^{(i)}\}\subset\Delta_{2}$. The ground-truth weight $\lambda^{*}$ corresponding to $\mathbb{Y}$ is marked by a red cross. The yellow and green crosses indicate the estimates of $\lambda^{*}$ produced by Algorithms 2 and 4, respectively. The three panels correspond to three independent synthesis runs used to generate the barycenters $\{\mathbb{Y}_{\lambda^{(i)}}\}$ over the simplex grid, incorporating the technical safeguards introduced to mitigate large run-to-run variability.
Figure 20. Heatmaps illustrating the landscape of $\mathbf{J}$ over the first two components of the simplex $\Delta_{2}$. For the same three fixed templates and the same target $\mathbb{Y}$ outside their GW barycenter space, we evaluate $\mathbf{J}(\lambda^{(i)})$ on a grid $\{\lambda^{(i)}\}\subset\Delta_{2}$. The yellow and green crosses indicate the estimates of GW $\lambda$-coordinates corresponding to $\mathbb{Y}$ produced by Algorithms 2 and 4, respectively. The three panels correspond to three independent synthesis runs used to generate barycenters $\{\mathbb{Y}_{\lambda^{(i)}}\}$ over the simplex grid, incorporating the technical safeguards introduced to mitigate large run-to-run variability.
E.1. Behavior Under Permutation of the Proposed Analysis Algorithms 2 and 4
From the theoretical perspective, both of our analysis methods are permutation invariant: if we relabel the nodes of a network, the $\lambda$-estimation does not change.
  • In Section 3.2, we address permutation/relabeling invariance in Remark 3.8: The formulation of the analysis problem considered there (see problem (13)), in which the GW barycenter model is taken over shapes in $\mathrm{Bary}_{M,\mathrm{q}}$ (Definition 7), together with its surrogate (16), are permutation/relabeling invariant. In both problems (13), (16) we take the divergence $d_{a}(\cdot,\cdot)$ to be the Frobenius norm on the space of $M\times M$ matrices (which allows us to develop Algorithm 2). Precisely, for fixed templates $\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]}$, given a network $(\mathbf{Y},\mathrm{q})$ and using the notation from Section 3.2, the landscape of the functional
    (65) $\Delta_{S-1}\ni\lambda\longmapsto\left\|\mathbf{Y}-\sum_{s\in[S]}\lambda_{s}\,F(\mathbf{Y},s)\right\|_{\mathrm{Frob}}=\lambda^{T}\mathcal{Q}\lambda$
    is invariant under relabeling the nodes of $(\mathbf{Y},\mathrm{q})$.
  • In Section 4.4, in formulating the analysis problem that leads to Algorithm 4, we work with the weak barycenter model $\widetilde{\mathrm{Bary}}$. Moreover, we consider equivalence classes (denoted as $\mathbb{Y}]$ for weakly isomorphic networks) which already incorporate permutation/relabeling invariance: that is, if the network $\mathbb{Y}$ is a representative of the class $\mathbb{Y}]$, and we relabel the nodes of such network creating a new one $\mathbb{Y}^{\prime}$, then $\mathbb{Y}]=[\mathbb{Y}^{\prime}]$. Precisely, for fixed templates $\{[\mathbb{X}^{s}]\}_{s\in[S]}$, given a network $\mathbb{Y}=(\mathbf{Y},\mathrm{q})$ and using the notation from Section 4.4, the landscape of the functional
    (66) $\Delta_{S-1}\ni\lambda\longmapsto\mathrm{tr}_{\mathrm{q}_{b}}\left(\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}-\mathbf{Y}_{b}\right)^{T}\left(\sum_{r=1}^{S}\lambda_{r}\mathbf{X}^{r}_{b}-\mathbf{Y}_{b}\right)\right)=\lambda^{T}\mathcal{A}\lambda$
    is invariant under relabeling the nodes of $\mathbb{Y}$.
In this section we demonstrate this invariance through experimental visualizations. Please see Fig. 21. For all the plots, we consider $3$ templates as point clouds selected as describe above in the main part of E. Equal sized point clouds, with uniform weights across nodes, where used in all these illustrative experiments.
a Inside Barycenter Space – FP landscape.
b Outside Barycenter Space - FP landscape.
c Inside Barycenter Space – BU landscape.
d Outside Barycenter Space – BU landscape.
Figure 21. Permutation/relabeling invariance examples for fixed-point (FP) analysis Algorithm 2 and blow-up (BU) analysis Algorithm 4. We only visualize the first two coordinates of the simplex $\Delta_{2}$. For each subplot (a)–(d), we consider a point cloud target $(\mathbf{Y},\mathrm{q})$ and two permutations (node relabeling) of it. On the left ((a) and (c)), the two input targets are synthetically generated as a GW barycenter (the red cross indicates the true =(0.3,0.4,0.3)=(0.3,0.4,0.3)). On the right ((b) and (d)), two targets are considered. Top panels (a) and (b) depict the heatmap of the landscape (65) when varying $\lambda\in\Delta_{2}$. Bottom panels (c) and (d) depict the heatmap of the landscape (66) when varying weights in $\Delta_{2}$. For all subplots (a)–(d), $\lambda$-estimation through Algorithm 2 is marked as a yellow cross, and $\lambda$-estimation via Algorithm 4 as a green cross in all plots. In each of the subfigures (a)-(d), the three corresponding landscapes are almost identical, demonstrating the invariance.
Appendix F Wasserstein Barycenters and Gromov-Wasserstein Barycenters
@@
(67) $\mathcal{P}_{2}(\mathbb{R}^{d})\ni\nu\longmapsto\frac{1}{2}\sum_{s=1}^{S}\lambda_{s}\,W^{2}_{2}(\nu,\mu^{s}).$
@@
(69) $\sum_{s=1}^{S}\lambda_{s}(Id-T^{\nu_{\lambda}\to\mu^{s}})=0\quad\text{almost everywhere w.r.t. }\nu_{\lambda}.$
Moreover, it can be proven that under certain assumptions critical points (or Karcher means), that is, measures $\nu_{\lambda}$ satisfying (69) and Wasserstein barycenters (i.e., minimizer of the functional described by (67)) coincide (see [72] and [46, Thm. 3.1.15]). As a consequence, to find such a Wasserstein barycenter $\nu_{\lambda}$, one can adopt the following fixed-point scheme based on (Wasserstein) gradient descent on the variance functional (67):
@@
In the context of this paper, that is, under the Gromov-Wasserstein (GW) structure, the objective functional $G_{\lambda}$ (see (30)) plays the role analogous to that of (67) in the Wasserstein setting. Its gradient can be computed using the blow-up technique (see Proposition 4.3), and it serves as the counterpart to the Fréchet derivative in (68). In [15], the authors propose an iterative Fréchet mean algorithm based on the work of Pennec [48], interpreting it as a gradient descent method for minimizing the Fréchet functional $G_{\lambda}$ over the space of finite networks. This approach parallels the procedure described in the Wasserstein setting but is adapted to the GW framework, following Pennec’s ideas on averaging in finite-dimensional Riemannian manifolds. However, to the best of our knowledge, there are no theoretical guarantees in the literature that critical points of $G_{\lambda}$ are minimizers of (2), that is, true GW barycenters. Our Lemma 4.5 represents a step in this direction.